OpenStudy (anonymous):

Factor 5x^4 - 405m^4

OpenStudy (zzr0ck3r):

factor out the 5 5(x^4-81m^4) = 5((x^2)^2 - (9m^2)^2) now use the fact that a^2-b^2 = (a-b)*(a+b)

OpenStudy (zzr0ck3r):

let me know what you get @Geodripp

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (zzr0ck3r):

then you will use that fact again

OpenStudy (anonymous):

Okay so I think I got it. -5(3m-x)(3m+x)(9m^2-x^2)